# SAGI: Semantically Aligned and Uncertainty Guided AI Image Inpainting

### v2

@@ 3.3 SAGI-D
Source of authentic data.
To ensure diversity and robustness, we leverage datasets spanning multiple domains: general object detection, high-resolution photography, and large-scale segmentation. Specifically, we utilize three publicly available datasets for authentic image sources: (1) MS-COCO [38], which provides images with captions and object masks across 80 categories (2) RAISE [18], a high-resolution dataset of 8,156 uncompressed RAW images designed for forgery detection evaluation (3) OpenImages [5], which offers extensive object segmentation data with over 2.7 million segmentations across 350 categories.
@@
For SAOR, we use dataset-provided masks and captions for COCO and OpenImages, while for RAISE, we use OneFormer [23] as the segmentation model $\Phi_{\text{seg}}$ and BLIP-2 [36] as the captioning model $\Psi_{\text{cap}}$. For prompt generation, we use ChatGPT 3.5 [50] (for COCO/RAISE) and Claude Sonnet 3.5 [3] (for OpenImages) as the language model $\Theta_{\text{llm}}$. The prompt engineering methodology is detailed in the supplementary material.
@@
For UGDA, we use GPT-4o [51] as the vision-language model $\Omega_{\text{vlm}}$, chosen for its effectiveness in synthetic image detection [80]. Empirically, we observed that QAlign effectively filters out low-quality images, so we applied UGDA to approximately half of the test inpainted images, selecting those with the highest QAlign scores. The complete prompt engineering methodology is detailed in the supplementary material.
@@
Diverse inpainting models are crucial for creating robust detection datasets and high-quality images, capturing varied artifacts, styles, and visual traits. Some images underwent a second inpainting round with different masks and prompts, simulating real-life multi-edit scenarios. We use five pipelines: HD-Painter [44], BrushNet [25], PowerPaint [93], ControlNet [86], and Inpaint-Anything [83], including its Remove-Anything variant for object removal. These pipelines support eight models, primarily based on Stable Diffusion [59], except Remove-Anything, which uses LaMa [65]. Due to memory limits, images were resized to a maximum dimension of 2048 pixels, except for Inpaint-Anything, which preserved original dimensions via cropping and resizing. One-sixth of the images underwent an additional inpainting round.
@@
We categorize inpainted images based on how models handle unmasked regions: if the unmasked region is preserved, we refer to them as Spliced (SP) images; if regenerated, as Fully Regenerated (FR) images. Inpaint-Anything preserves unmasked regions through copy-paste (SP), while ControlNet regenerates the full image (FR). BrushNet, PowerPaint, and HD-Painter can produce both SP and FR images depending on post-processing settings (e.g., blending or upscaling). Remove-Anything, based on LaMa architecture, inherently preserves unmasked regions, thus producing SP images. This diversity in processing approaches contributes to a more comprehensive dataset, as FR images are typically more challenging to detect than SP images [66].
@@
Our dataset evaluates both in-domain performance and generalization to new data. For in-domain evaluation, we use COCO (60,000 training and nearly all 5,000 validation images) and RAISE (7,735 images processed with $\Phi_{\text{seg}}$, yielding 25,674 image-mask-model combinations). For generalization, we create an out-of-domain split using OpenImages (6,000 test images), employing a different language model $\Theta_{\text{llm}}$ (Claude) compared to COCO and RAISE (ChatGPT). This tests performance on new data and a different LLM. Detailed statistics are in the supplementary material.
@@ 4 Experimental Evaluation
Our experimental evaluation focuses on two main directions: (1) assessing the quality, aesthetics, and realism of inpainted images, and (2) evaluating the effectiveness of SAGI-D as a benchmark for inpainting detection.
@@ 4.1 Image Quality and Realism Evaluation
We conduct two ablations to validate our design choices. First, we evaluate SAOR’s use of language models by comparing three prompting approaches: (1) using just object labels, (2) combining object labels with image captions, and (3) feeding both labels and captions to an LLM to generate enhanced prompts. Second, we assess UGDA’s effectiveness through a human study with 42 participants on 1,000 images, comparing human perception against model performance on images classified as deceiving or undeceiving.
@@
Table 1: Comparison of metrics across inpainting models with different prompt types. The SAOR columns indicate which the prompting approach: object labels (Obj), image captions (Cap), and LLM-generated prompts (LLM). Bold values indicate better performance per model-metric pair. Pipelines: BN (BrushNet), CN (ControlNet), HDP (HD-Painter), IA (Inpaint-Anything), PPt (PowerPaint). Metrics: AS (Aesthetics Score), CS Ae (Clip Aesthetics), QA Ae (QAlign Aesthetics), CS Qlt (Clip Quality), QA Qlt (QAlign Quality). Notably, LLM prompts substantially enhance aesthetics, while maintaining performance even in quality metrics like CS Qlt, that are typically less sensitive to semantic variations.
@@
Dec Und Diff Dec Und Diff
UGDA 35.2 73.7 +38.5 39.9 12.7 +26.8
Table 2: Comparison of human accuracy and mean IoU on 500 inpainted images, separated into two groups for each metric. For AS, QA Qlt, CS, and QA Ae, groups are formed by ranking images and splitting them into the top 50% and bottom 50% based on their metric scores; for UGDA, groups are based on its assessment (deceiving/undeceiving). The difference column shows the absolute difference in accuracy between the two groups. UGDA can clearly seperate deceiving from undeceiving images.
@@
Table 3: Comparison of Inpainting Dataset Characteristics. Our dataset surpasses existing ones in scale (number of images), diversity (source datasets, models, pipelines). Resolution varies based on source dataset. AI: AI-Generated Content, OR: Object Removal.
To compare LLM-generated prompts with object labels or labels combined with captions, we evaluated 900 original images across five inpainting models, generating 4,500 images per prompt type (LLM, object labels, and caption prompts). We compared aesthetics and quality metrics, including CLIP Similarity [55] for aesthetics (CS Ae) [22] and quality (CS Qlt) [74] , QAlign [77] for quality (QA Qlt) and aesthetics (QA Ae), and Aesthetic Score (AS) [62]. Tab. 1 demonstrates that LLM prompts consistently outperform both object labels and object labels along with caption prompts across most metrics and models, enhancing both image quality and aesthetics. For instance, all models achieve higher aesthetic metrics (QA Ae, CS Ae, AS) with LLM prompts. BrushNet and ControlNet excelled in all metrics, while HD-Painter, PowerPaint, and Inpaint-Anything showed slight variations when using just object or caption labels in QA Qlt and CS Qlt. Quality improvements were marginal compared to aesthetic gains, as quality metrics focus on technical aspects less influenced by prompt content. On average, LLM prompts achieved higher scores across all metrics except QA Qlt when compared to captions. These results highlight the richer semantic guidance of LLM prompts, enhancing inpainting quality and aesthetics.
@@
To validate UGDA’s effectiveness and establish a human baseline, we conducted a comprehensive user study comparing human perception with model detection capabilities (see supplementary material for additional results). We created a balanced evaluation set of 1,000 images: 250 classified as deceiving by UGDA, 250 as undeceiving, and 500 unaltered images as control. Images were randomly selected with constraints to avoid redundancy: no overlap between authentic images and originals of inpainted images, and no multiple inpainted versions from the same source. The study involved 42 participants evaluating batches of 20 images, with each image receiving 3-5 independent assessments. Detailed participant statistics are provided in the supplementary material.
Table 2 compares human accuracy and mean IoU on the 500 inpainted images, split into two groups for each metric. For AS, QA Qlt, CS, and QA Ae, groups are based on metric rankings (top 50% and bottom 50%); for UGDA, groups are based on its assessment (deceiving/undeceiving). The results show that existing quality and aesthetics metrics (AS, QA Qlt, CS Qlt, CS Ae, QA Ae) struggle to distinguish between convincing and unconvincing images, with minimal differences in accuracy in the two groups (0.4 to 12.9) and IoU (0.2 to 8.3). In contrast, UGDA achieves a much larger difference (38.5 in accuracy and 26.8 in IoU), demonstrating its ability to effectively identify realistic inpaintings. This highlights the need for models like UGDA for distinguishing realistic from unrealistic inpaintings.
@@
To the best of our knowledge, SAGI-D is the largest collection of AI-generated inpainted images (see supplementary material for additional metrics and qualitative analysis). Tab. 3 compares SAGI-D with existing datasets. With 95,839 inpainted images from 77,900 originals, it surpasses TGIF (74,976 from 3,124) and CocoGlide (512 from 512) in scale. Unlike single-source datasets, SAGI-D integrates COCO, RAISE, and OpenImages, enhancing diversity. It employs eight inpainting models across five pipelines, enabling complex manipulations like double inpainting, and supports resolutions up to 2048p, exceeding TGIF’s 1024p and CocoGlide’s 256p. SAGI-D uniquely includes a human benchmark subset and an out-of-domain test set, making it valuable for evaluating inpainting detection under diverse conditions. It outperforms existing datasets in aesthetic and quality metrics (CS, QA Qlt, QA AE, AS), demonstrating higher perceptual alignment with human judgment and improved visual fidelity. By combining AI-generated content (AIGC) and object removal (OR) techniques, along with human-evaluated and out-of-domain subsets, SAGI-D offers a comprehensive benchmark for inpainting detection across varied manipulation scenarios. This positions SAGI-D as a superior resource for advancing detection models and sets a new standard for benchmarking in the field.
@@ 4.2 Inpainting Detection Benchmark
To establish a comprehensive benchmark for the presented SAGI-D dataset, we evaluate the performance of several state-of-the-art image manipulation detection models.
@@
Given an RGB image $x^{rgb}\in\mathbb{R}^{(H\times W\times 3)}$, the inpainting detection model aims to predict either a pixel-level inpainting localization mask $\hat{y}^{loc}\in(0,1)^{(H\times W\times 1)}$ and/or an image-level inpainting detection probability $\hat{y}^{det}\in(0,1)$. The former will be referred to as a localization problem, and the latter as a detection problem.
@@
We evaluated four inpainting detectors: PSCC [39], CAT-Net [34], TruFor [20], and MMFusion [71]. CAT-Net provides only pixel-level masks, so we used the maximum probability from these masks for image-level detection. The other models output both pixel localization and image-level detection probabilities.
@@
We evaluated the pretrained models, the versions retrained on the SAGI-D and TGIF following the training protocol on the original papers.
@@
We retrained all models from scratch, following the training protocol outlined in their original papers. CAT-Net was trained on an NVIDIA A100 GPU, while PSCC, TruFor, and MMFusion were trained on an NVIDIA RTX 4090.
@@
We evaluated model performance at image and pixel levels using distinct metrics. With the positive class representing inpainted regions, for image-level detection, we use accuracy, and for pixel level we measure Intersection over Union (IoU), with a threshold of $0.5$. To compare models across datasets, we used the threshold-agnostic Area Under the Curve (AUC) metric. For localization AUC, we resized and flattened localization maps and ground truths into vectors for ROC computation (see supplementary for additional results, qualitative comparisons, and human performance analysis).

### v3

@@ 3.3 SAGI-D
Sources of authentic images.
To ensure diversity and robustness, we leverage datasets spanning multiple domains: general object detection, high-resolution photography, and large-scale segmentation. Specifically, we utilize three publicly available datasets for authentic image sources: (1) MS-COCO [40], which provides images with captions and object masks across 80 categories, (2) RAISE [22], a high-resolution dataset of 8,156 uncompressed RAW images designed for forgery detection evaluation and (3) OpenImages [35], which offers extensive object segmentation data with over 2.7 million segmentations across 350 categories.
@@
We use dataset-provided masks and captions for MS-COCO and OpenImages, while for RAISE, we use OneFormer [27] as the segmentation model $\Phi_{\text{seg}}$ and BLIP-2 [38] as the captioning model $\Psi_{\text{cap}}$. To generate prompts, we employ as the language model $\Theta_{\text{llm}}$ ChatGPT 3.5 [53] for MS-COCO and RAISE as well as Claude Sonnet 3.5 [4] for OpenImages. The prompt engineering methodology is detailed in the supplementary material.
@@
We use GPT-4o [54] as the VLM $\Omega_{\text{vlm}}$, chosen for its effectiveness in synthetic image detection [84] and for achieving the best performance in our respective ablation studies presented in Tab. 3. Empirically, we observed that QAlign effectively filters out low-quality images, so we applied UGDA to approximately half of the test inpainted images, selecting those with the highest QAlign scores. The complete prompt engineering methodology is detailed in the supplementary material.
@@
Diverse inpainting models are crucial for creating robust detection datasets and high-quality images, capturing varied artifacts, styles, and visual traits. We use five pipelines: HD-Painter [46], BrushNet [29], PowerPaint [97], ControlNet [90], and Inpaint-Anything [87], including its Remove-Anything variant for object removal, with their original padding and normalization strategies maintained. These pipelines support eight models, primarily based on Stable Diffusion [62], except Remove-Anything, which uses LaMa [68]. One-sixth of the images underwent a second inpainting round with different masks and prompts, simulating real-life multi-edit scenarios. Also, due to memory limits, images were resized to a maximum dimension of 2048 pixels, except for Inpaint-Anything, that preserves original dimensions via cropping and resizing.
@@
We categorize inpainted images based on how models handle unmasked regions: if the unmasked region is preserved, we refer to them as Spliced (SP) images; if regenerated, as Fully Regenerated (FR) images. Inpaint-Anything preserves unmasked regions through copy-paste (SP), while ControlNet regenerates the full image (FR). BrushNet, PowerPaint, and HD-Painter can produce both SP and FR images depending on post-processing settings (e.g., blending or upscaling). Remove-Anything inherently preserves unmasked regions, thus producing SP images. This diversity in processing approaches contributes to a more comprehensive dataset, as FR images are typically more challenging to detect than SP images [69].
@@
Our dataset considers both in-domain performance as well as generalization to new data. Regarding in-domain evaluation, we use MS-COCO (60,000 training and nearly all 5,000 validation images) and RAISE (7,735 images processed with $\Phi_{\text{seg}}$, yielding 25,674 image-mask-model combinations). To test generalization across different image sources and LLMs, we create an out-of-domain split from OpenImages (6k test images), by also using a different language model $\Theta_{\text{llm}}$, i.e. Claude versus ChatGPT for COCO and RAISE. Detailed statistics are provided in the supplementary material.
@@ 4 Experimental Evaluation
Our experimental evaluation focuses on two main directions: (1) assessing the quality, aesthetics, and realism of images inpainted with SAGI, and (2) evaluating the effectiveness of SAGI-D as a benchmark for inpainting detection.
@@ 4.1 Image Quality and Realism Evaluation
We conduct two ablations to validate our design choices. First, we evaluate SAOR’s use of language models by comparing three prompting approaches: (1) using just object labels, (2) combining object labels with image captions, and (3) feeding both labels and captions to an LLM to generate enhanced prompts. Second, we assess UGDA’s effectiveness through a human study with 42 participants on 1,000 images, comparing human perception against model performance on images classified as deceiving or non-deceiving. We also ablate the VLM employed in UGDA, by evaluating the ability of several popular VLMs to align with human perception of image realism.
@@
Table 1: Comparison of metrics across inpainting models with different prompt types. The SAOR columns indicate the employed prompting method: object labels (Obj), image captions (Cap), and LLM-generated prompts (LLM). Bold values indicate better performance per model-metric pair. Pipelines: BN (BrushNet), CN (ControlNet), HDP (HD-Painter), IA (Inpaint-Anything), PPt (PowerPaint). Metrics: AS (Aesthetics Score), CS Ae (Clip Aesthetics), QA Ae (QAlign Aesthetics), CS Qlt (Clip Quality), QA Qlt (QAlign Quality). LLM prompts notably enhance aesthetics, while maintaining performance even in quality metrics like CS Qlt, that are typically less sensitive to semantic variations.
@@
Dec Non-Dec Diff Dec Non-Dec Diff
UGDA 35.2 73.7 +38.5 12.7 39.9 +26.8
Table 2: Comparison of human accuracy and mean IoU on 500 inpainted images, separated into two groups for each metric. For AS, QA Qlt, CS, and QA Ae, groups are formed by ranking images and splitting them into the top 50% and bottom 50% based on their metric scores; for UGDA, groups are based on its assessment (deceiving/non-deceiving). UGDA can clearly separate deceiving from non-deceiving images.
VLM / Human Eval. Acc $\downarrow$ IoU $\downarrow$
w/o UGDA 54.954.954.9 26.426.426.4
Gemma-3-27b [72] $45.4^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+9.5}}$ $20.0^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+6.4}}$
Mistral-S-3.1-24B [2] $45.2^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+9.7}}$ $20.4^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+6.0}}$
Qwen2.5-VL-7B [6] $45.2^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+9.7}}$ $21.0^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+5.4}}$
Claude-3-7-sonnet [5] $43.4^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+11.5}}$ $17.8^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+8.6}}$
Gemini-2.5-flash [19] $38.6^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+16.3}}$ $16.5^{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+9.9}}$
GPT-4o [54] $\textbf{35.2}^{\textbf{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+19.7}}}$ $\textbf{12.7}^{\textbf{{\color[rgb]{0,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,0}+13.7}}}$
Table 3: VLM ablation for UGDA. Human accuracy and mean IoU for distinguishing inpainted from real images after filtering with different VLMs. Lower values indicate more deceiving inpaintings. All VLMs improve over the unfiltered baseline, with GPT-4o achieving the strongest alignment with human perception.
@@
Table 4: Comparison of inpainting datasets characteristics. Our dataset surpasses existing ones in scale (number of images), diversity (source datasets, models, pipelines). Resolution varies based on source dataset. AI: AI-Generated Content, OR: Object Removal.
To compare LLM-generated prompts with object labels or labels and captions, we evaluated 900 original images across five inpainting models, generating 4,500 images per prompt type. We compared aesthetics and quality metrics, including CLIP Similarity [58] for aesthetics (CS Ae) [26] and quality (CS Qlt) [78] , QAlign [81] for quality (QA Qlt) and aesthetics (QA Ae), and Aesthetic Score (AS) [65]. Tab. 1 shows that LLM prompts consistently outperform object labels and caption prompts across most metrics and models, enhancing both image quality and aesthetics. All models achieve higher aesthetic metrics with LLM prompts. BrushNet and ControlNet excelled across all metrics, while other models showed slight variations in quality metrics. Quality improvements were marginal compared to aesthetic gains, as quality metrics primarily focus on technical aspects of the image which are less influenced by prompt’s content. Overall, these results highlight the richer semantic guidance of LLM prompts, significantly enhancing inpainting aesthetics, while also improving some technical aspects of the generated visual content.
@@
We validated UGDA through a user study with 42 participants on 1,000 images: 250 inpainted ones classified by UGDA as deceiving, 250 as non-deceiving and 500 authentic. Images were selected to avoid redundancy, i.e. no authentic-inpainted overlap and no multiple inpainted versions per authentic source image. Participants evaluated batches of 20 images, with each image receiving 3-5 independent assessments. Detailed participant statistics are provided in the supplementary material.
Tab. 2 compares human accuracy and mean IoU on 500 inpainted images, split into two groups per metric. For AS, QA Qlt, CS, and QA Ae, splits are based on top/bottom 50% rankings; for UGDA, on its deceiving/non-deceiving assessment. Existing quality and aesthetic metrics (AS, QA Qlt, CS, QA Ae) show minimal differences among groups (accuracy: 0.4–12.9; IoU: 0.2–8.3), indicating limited correlation with human perception of realism. In contrast, UGDA yields much larger gaps (accuracy: 38.5; IoU: 26.8), effectively identifying realistic inpaintings. This underscores the need for uncertainty guidance to capture realism.
To further validate the efficacy and model-agnostic nature of UGDA, we compare the performance of several VLMs. Tab. 3 compares human ability to distinguish inpainted from real images, both in terms of accuracy and mean IoU, after non-deceiving samples have been filtered out using our methodology. Results demonstrate that our method enables all tested VLMs to better align with human preference and discard non-realistic inpaintings, as evidenced by the positive differences over a baseline without UGDA-based filtering. As GPT-4o shows the strongest alignment with human perception, we employ it as the primary model in our benchmark creation.
@@
To the best of our knowledge, SAGI-D is the largest collection of AI-generated inpainted images. Tab. 4 compares SAGI-D with existing datasets. With 95,839 inpainted images from 77,900 originals, it surpasses TGIF (74,976 from 3,124) and CocoGlide (512 from 512) in scale. Unlike single-source datasets, SAGI-D integrates COCO, RAISE, and OpenImages, enhancing diversity. It employs eight inpainting models across five pipelines, considers advanced manipulations such as double inpainting, AI generated content, and object removal, and allows resolutions up to 2048p—exceeding TGIF’s 1024p and CocoGlide’s 256p. SAGI-D uniquely includes human benchmark and out-of-domain test subsets, making it valuable for evaluating detection under diverse conditions. It outperforms existing datasets in aesthetic and quality metrics, demonstrating superior perceptual alignment and visual fidelity. This positions SAGI-D as a superior resource for advancing detection models and sets a new standard for benchmarking in the field.
Failure Cases
While our framework demonstrates strong overall performance, certain limitations remain. A representative failure case for SAOR occurs when, despite the context provided, the LLM generates semantically incoherent prompts. For example, in the first row of Fig. 4, a hot air balloon is suggested as a replacement for a train object. For UGDA, the VLM can exhibit inconsistent behavior by fixating on specific details while ignoring obvious artifacts. As shown in the second row of Fig. 4, in the first permutation, the VLM insists on “couch duplication” and fails to recognize clear blurring artifacts, whereas in the second permutation, it correctly identifies the same blur. These cases highlight the dependency of our approach on the underlying capabilities and biases of foundation models. More examples are discussed in the supplementary material.
@@ 4.2 Inpainting Detection Benchmark
To establish a comprehensive benchmark for the presented SAGI-D dataset, we evaluate the performance of several state-of-the-art image forgery detection models.
@@
Given an RGB image $x^{rgb}\in\mathbb{R}^{(H\times W\times 3)}$, the inpainting detection model aims to predict a pixel-level inpainting localization mask $\hat{y}^{loc}\in(0,1)^{(H\times W\times 1)}$ and/or an image-level inpainting detection probability $\hat{y}^{det}\in(0,1)$. The former will be referred to as the localization task, and the latter as the detection task.
@@
We evaluated four inpainting detectors: PSCC-Net [41], CAT-Net [36], TruFor [24], and MMFusion [75]. CAT-Net provides only pixel-level masks, so we used the maximum probability from these masks for image-level detection. The other models output both pixel localization and image-level detection probabilities.
@@
We evaluated the pretrained models, the versions retrained on the SAGI-D and on TGIF [48] following the training protocol on the original papers.
@@
We retrained all models from scratch, following the training protocol outlined in their original papers. CAT-Net was trained on an NVIDIA A100 GPU, while PSCC-Net, TruFor, and MMFusion were trained on an NVIDIA RTX 4090.
@@
We evaluated model performance at image and pixel levels using distinct metrics. With the positive class representing inpainted regions, for image-level detection, we use accuracy, and for pixel level we measure Intersection over Union (IoU), with a threshold of 0.50.50.5. To compare models across datasets, we used the threshold-agnostic Area Under the Curve (AUC) metric. For localization AUC, we resized and flattened localization maps and ground truths into vectors for ROC computation (see supplementary for additional results, qualitative comparisons, and human performance analysis).